If 40.0dmcube of nitrogen is collected over water of 22 °c when atmospheric pressure Is 727 Torr,what is the volume of dry nitrogen at standard temperature and pressure, assuming ideal behavior?


1
Expert's answer
2022-03-17T15:01:20-0400

We use the equation: P1V1/T1 = P2V2/T2

Convert the units:

T = 220C = 295.15 K – initial temperature

At STP P = 760 Torr

At 220C vapor pressure of water is 19.5 Torr

We subtract it from general pressure:

760 – 19.5 = 740.5 Torr

T = 298 K – standard temperature

727*40.0/295 = 740.5*V2/298

98.58 = 2.48 V2

V2 = 39.75 dm3